OpenStudy (anonymous):

What is a polynomial function in standard form with zeros 1, 2, -3, and -1?

OpenStudy (anonymous):

A) g(x)=x^4+x^3-7x^2-x+6 B )g(x)=x^4+x^3+7x^2-x+6 C) g(x)=x^4+x^3-7x^2-x-6 D) g(x)=x^4-x^3-7x^2-x+6

OpenStudy (anonymous):

Hint: A quartic with roots \[\alpha, \beta, \gamma, \delta\]can be written as: \[(x-\alpha)(x-\beta)(x-\gamma)(x-\delta) = 0\]